40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6967     Entry time: Thu Jul 12 13:27:16 2012
Author: yuta 
Type: Update 
Category: SUS 
Subject: measured OLTFs of PRMI optics' SUSPOS 

I measured openloop transfer functions of PRM/BS/ITMX/ITMY SUSPOS servo.
They look great. They all have ~50 deg phase margin and damps only POS resonance.
PRMSUSPOSOLTF.pngBSSUSPOSOLTF.pngITMXSUSPOSOLTF.pngITMYSUSPOSOLTF.png

ELOG V3.1.3-